# A motivic Fundamental Lemma

### v3

@@ 1. Introduction
[…]
@@
with $[\gamma^{\prime}]$ running over the set of $G(F)$-conjugacy classes within the stable conjugacy class of $\gamma_{G}$. When $\kappa$ is the trivial character one gets the stable orbital integral $O^{\mathrm{stab}}_{\gamma_{G}}$. Thus, the problem of stabilization of orbital integrals can be reduced to expressing a $\kappa$-orbital integral for $G$ in terms of a stable orbital integral on some different unramified reductive group $H$ over $F$.
@@
The Hitchin fibration is endowed with an action of a Picard stack ${\mathcal{P}}_{G}\to\operatorname{\mathsf{A}}$. The anisotropic locus $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$ in $\operatorname{\mathsf{A}}$ is the locus of points $a$ such that every Higgs bundle $(E,\theta)\in\chi^{-1}(a)$ is generically anisotropic in the sense that for any trivialization of $E$ and $D$ over the generic point of $X$, the Higgs field $\theta$ is anisotropic as an element of $\mathfrak{g}(k(X))$. When $a$ belongs to $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, the group $\pi_{0}({\mathcal{P}}_{G,a})$ is finite. Fix $a\in\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, $H$ an endoscopic group of $G$, and $\kappa:\pi_{0}({\mathcal{P}}_{G,a})\to\overline{\mathbb{Q}_{\ell}}^{\times}$ the corresponding character. For technical reasons one needs to work on some étale open covering $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}\to\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, and we denote by ~$\widetilde{\mathcal{M}}_{G}$, etc, the corresponding pullback. The locus $\widetilde{\operatorname{\mathsf{A}}}_{H}^{{\mathrm{ani}}}$ relative to $H$ is contained in $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}$.
@@
#$\#^{\kappa}\,\widetilde{{\mathcal{M}}}_{G,a}(k)=q^{r_{G}^{H}(D)}\#^{\mathrm{\mathrm{stab}}}\,\widetilde{{\mathcal{M}}}_{H,a}(k).$
Here #$\#^{\kappa}$ denotes the trace of Frobenius on the $\kappa$-isotypical component of the cohomology with compact supports, #$\#^{\mathrm{stab}}$ denotes #$\#^{\kappa}$ when $\kappa$ is the trivial character, $q$ is the cardinality of $k$ and $r_{G}^{H}(D)$ denotes $1/2(\mathrm{dim}(\widetilde{\mathcal{M}}_{G})-\mathrm{dim}(\widetilde{\mathcal{M}}_{H}))$.
@@ 2.1. Pseudo-finite fields
Let $\mathrm{\bf K}(\mathrm{Var}_{S})$ be the Grothendieck group of varieties over $S$ and $\mathcal{M}_{S}$ its localization by the class $\mathbb{L}$ of the affine line over $S$. Ivorra and Sebag prove in [30, Lemma 2.1] the existence of a unique ring morphism
@@
By additivity of $p_{!}$, as in the proof of [30, Lemma 2.1], this extends uniquely to a group morphism
@@
Property $(1)$ now follows from the standard fact that $\underline{\mathrm{Hom}}(V_{\alpha},M)^{G}\otimes V_{\alpha}$ is isomorphic to $M^{\alpha}$, see [11, 6.1] for details. Property $(2)$ follows from the fact that $[M]=\sum_{\alpha\in G^{*}}[M^{\alpha}]$. We finally define $\chi_{S}(-,\alpha)$ for a virtual character $\alpha$ in the unique way such that $(3)$ holds. ∎
@@ 2.5. Denef-Pas language and definable subassignments
Fix a field $k$ of characteristic zero. The theory takes place in the Denef-Pas 3-sorted language $\mathcal{L}_{\mathrm{DP},k}$ in the sense of first order logic. It has one sort for the valued field with the ring language, one sort for the residue field with ring language, and one sort for the value group, with Presburger arithmetic language (0,1,+$\leq$, $(\equiv_{n})_{n\in\operatorname{\mathsf{N}}})$. It also has two unary function symbols, $\operatorname{ord}$ from the valued field sort to the value group, interpreted as the valuation, and $\operatorname{ac}$ from the valued field sort to the residue field, interpreted as an angular component. It also has constant symbols for elements of $k\mathopen{(\!(}t\mathopen{)\!)}$.
@@
Let $*_{k}$ the final object in $\mathrm{GDef}_{\mathrm{Psf},k}$, whose value at every pseudo-finite extension of $k$ is the one point set. By definition
@@ 2.6. Evaluation of functions
Since $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k,x}$ is an expansion of $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k}$, by [9, Proposition 16.1.1], every $\psi\in\mathcal{C}_{\mathrm{Psf},k}(S)$ induces a unique function in $\mathcal{C}_{\mathrm{Psf},k,x}(S)$, that we still denote abusively $\psi$. By cell decomposition, or more precisely [9, Theorem 7.2.1], there exists an $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k,x}$-definable subassignment $V_{x}\subseteq S$, such that $x\in V_{x}(K)$, a definable function $g\colon V_{x}\to*$ and $\psi_{x}\in\mathcal{C}_{\mathrm{Psf},k,x}(*)$ such that $g^{*}(\psi_{x})=\psi$. We define the evaluation of $\psi$ at $(x,K)$ to be $\psi(x)=\psi_{x}\in\mathcal{C}_{\mathrm{Psf},k,x}(*)=\mathcal{C}_{\mathrm{Psf},k(x)}(*_{k(x)})$
@@ 2.7. Definable varieties
We shall usually consider such ${\mathbf{X}}$ as a definable in ${\mathbb{P}}^{n}$ with parameters $\pi({\mathbf{X}})$. In particular, by “${\mathbf{X}}$ is of pure dimension $d$” we mean that the fibers ${\mathbf{X}}_{a}$ are points of varieties of pure dimension $d$.
@@
$\mathbf{U}_{1}=\left\{(x_{1},x_{2},a)\in{\mathbb{A}}^{2}\times{\mathbb{A}}\mid x_{1}^{2}-ax_{2}^{2}=1,a\text{ not a square }\right\}.$
@@ 2.8. Fubini theorem
We say that $S\in\mathrm{GDef}_{\mathrm{Psf},k}$ is of dimension $d$ if $\mathrm{Kdim}(S)=d$, in the sense of [9, Section 3]. Roughly speaking, this means that the valued field part of $S$ is of dimension $d$. Fix $S\in\mathrm{Def}_{\mathrm{Psf},k}$ of dimension $d$ and let $\varphi\in\mathcal{C}_{\mathrm{Psf}}(S)$. Let $\left\lvert\omega_{0}\right\rvert_{S}$ be the canonical volume form, in the sense of [9, 15.1]. Let $C_{\mathrm{Psf}}^{d}(S)$ the quotient of $\mathcal{C}_{\mathrm{Psf}}(S)$ by the ideal of constructible functions with support included in a definable of dimension $d-1$. If the class $[\varphi]\in C_{\mathrm{Psf}}^{d}(S)$ of $\varphi$ is integrable, then $\int_{S}[\varphi]\left\lvert\omega_{0}\right\rvert_{S}$ is defined in [9, 15.1] and belongs to $\mathcal{C}_{\mathrm{Psf}}(*_{k})$. Say that $\varphi\in\mathcal{C}_{\mathrm{Psf}}(S)$ is integrable if $[\varphi]$ is integrable, and set
@@
Let ${\mathbf{X}}$ be a smooth definable algebraic variety over $k\mathopen{(\!(}t\mathopen{)\!)}$ of pure dimension $d$. Denote by $\left\lvert\omega_{\mathbf{X}}\right\rvert$ the definable volume form on ${\mathbf{X}}$ determined by the algebraic differential form of degree $d$ on ${\mathbf{X}}_{a}(K\mathopen{(\!(}t\mathopen{)\!)})$ for every pseudo-finite field. Let $\varphi\in\mathcal{C}_{\mathrm{Psf}}({\mathbf{X}})$. We say that $\varphi\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable if $[\varphi]\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable and set
@@ 2.9. Constructible functions with values in motives
Let $S\in\mathrm{GDef}_{\mathrm{Psf},k}$ be of dimension $d$. Let $\mathcal{C}^{\leq k}_{\mathrm{Psf}}(S)$ be the subring of $\mathcal{C}_{\mathrm{Psf}}(S)$ consisting of constructible functions with support contained in a definable subassignment of dimension at most $k$, and $C^{\leq k}_{\mathrm{Psf}}(S)=\mathcal{C}^{\leq k}_{\mathrm{Psf}}(S)/\mathcal{C}^{\leq{k-1}}_{\mathrm{Psf}}(S)$. Let $C_{\mathrm{Psf}}(S)=\bigoplus_{k\geq 0}C^{k}_{\mathrm{Psf}}(S)$. Recall from [9, Section 10] that the integral is constructed by defining for each morphism $f\colon X\to Y$ in $\mathrm{Def}_{\mathrm{Psf},k}$ a group homomorphism $f_{!}\colon IC_{\mathrm{Psf}}(X)\to C_{\mathrm{Psf}}(Y)$, where $IC_{\mathrm{Psf}}(X)$ is the subgroup of $C_{\mathrm{Psf}}(X)$ consisting of integrable functions. The integral of $[\varphi]\in C_{\mathrm{Psf}}(X)$ is then defined as $f_{!}[\varphi]$, where $f\colon X\to*_{k}$ is the map to the final object. To prove Proposition 2.9.3, it is thus enough to prove that for every definable map $f\colon X\to Y$, $f_{!}\colon IC_{\mathrm{Psf}}(X)\to C_{\mathrm{Psf}}(Y)$ induces a map $IC_{\mathrm{mot}}(X)\to C_{\mathrm{mot}}(Y)$, where $IC_{\mathrm{mot}}(X)$ is the image of $IC_{\mathrm{Psf}}(X)$ by $\vartheta_{X}$.
@@
[…]
@@
Let ${\mathbf{X}}$ be a smooth definable algebraic $k\mathopen{(\!(}t\mathopen{)\!)}$-variety of pure dimension and $\left\lvert\omega_{\mathbf{X}}\right\rvert$ the associated volume form on ${\mathbf{X}}$. Fix $\varphi\in\mathcal{C}_{\mathrm{Psf}}({\mathbf{X}})$ such that $\varphi\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable. Using Proposition 2.9.3 on affine charts, we see that $\vartheta_{*_{k}}(\int_{{\mathbf{X}}}\varphi\left\lvert\omega_{\mathbf{X}}\right\rvert)$ depends only on $\vartheta_{{\mathbf{X}}}(\varphi)$. In this case, we say that $\vartheta_{{\mathbf{X}}}(\varphi)\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable and we set
$\int_{{\mathbf{X}}}^{\mathrm{mot}}\vartheta_{{\mathbf{X}}}(\varphi)\left\lvert\omega_{\mathbf{X}}\right\rvert=\vartheta_{*_{k}}\Bigl{(}\int_{{\mathbf{X}}}\varphi\left\lvert\omega_{\mathbf{X}}\right\rvert\Bigr{)}.$
@@ 2.13. Orbifold volume
A definable smooth Deligne-Mumford stack over $k\mathopen{[\![}t\mathopen{]\!]}$ is given by the following data: a definable groupoid $(V,A)$ such that $V/A$ is a definable algebraic variety ${\mathbf{M}}$ over $k\mathopen{[\![}t\mathopen{]\!]}$ together with a finite definable cover ${\mathbf{M}}=\cup{\mathbf{M}}_{i}$ such that the ${\mathbf{M}}_{i}$ form a Zariski cover of ${\mathbf{M}}$ into open pieces when specialized to a pseudo-finite field. We require as well that for all $i$, there is a smooth definable algebraic variety ${\mathbf{X}}_{i}$ and a definable finite algebraic group $\Gamma_{i}$ acting on ${\mathbf{X}}_{i}$ such that every orbit is contained in an affine subset, and such that the restriction of $(V,A)$ to ${\mathbf{M}}_{i}$ is identified with the quotient stack $[{\mathbf{X}}_{i}/\Gamma_{i}]\to{\mathbf{X}}_{i}/\Gamma_{i}\simeq{\mathbf{M}}_{i}$, and there is a dense open subset $\mathbf{U}_{i}\subseteq{\mathbf{M}}_{i}$ over which the action is free.
@@
After shrinking the ${\mathbf{X}}_{i}$ if necessary, we may assume that their canonical bundles are trivial. We fix non-vanishing top degree forms $\omega_{i}$ on $X_{i}$, which induce, possibly pluricanonical, forms $\omega_{i,\mathrm{orb}}$ on ${\mathbf{M}}^{\natural}_{i}$ as in [22, Lemma 2.8] and the measures induced by $\omega_{i,\mathrm{orb}}$ glue to define a definable form $\left\lvert\omega_{\mathrm{orb}}\right\rvert$ on ${\mathbf{M}}^{\natural}$. We shall compute $\int_{{\mathbf{M}}^{\natural}}\left\lvert\omega_{\mathrm{orb}}\right\rvert$ in terms of the twisted inertia stack of $\mathcal{M}_{K}$.
@@
${\mathcal{M}_{x}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x}$${M,}$
@@
Let ~$\widetilde{\mathcal{M}}_{x}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $\mathcal{M}_{x}$. By [22, Proposition 2.12], there is a finite totally ramified extension $L$ of $K\mathopen{(\!(}t\mathopen{)\!)}$ of degree $N$, with valuation ring $\mathcal{O}_{L}$ and residue field $k_{L}$, such that $[\operatorname{\mathsf{Spec}}(\mathcal{O}_{L})/\mu_{N}]\simeq\widetilde{\mathcal{M}}_{x}$ and the induced isomorphism
$[\operatorname{\mathsf{Spec}}(K)/\mu_{N}]=[\operatorname{\mathsf{Spec}}(k_{L})/\mu_{N}]\simeq\widetilde{\mathcal{M}}_{x,K}$
in the special fiber is independent of $L$ up to isomorphism. Hence we get a map $[\operatorname{\mathsf{Spec}}(K)/\mu_{N}]\to\widetilde{\mathcal{M}}_{x,K}\to\mathcal{M}_{x,K}\to\mathcal{M}_{K}$, which gives a point $e(x)\in I_{\hat{\mu}}M_{K}(K)=I_{\hat{\mu}}{\mathbf{M}}(K)$.
@@
Before starting the proof, recall from [22, Section 2.5] some facts about $\Gamma$-torsors over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$, with $K$ pseudo-finite. A $\Gamma$-torsor $T=\operatorname{\mathsf{Spec}}(\prod_{i}L_{i})$ over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$ is said to be unramified if each $L_{i}$ is unramified. It is strongly ramified if each $L_{i}$ is totally ramified.
@@
For any $T$, from the description of $\mathrm{H}^{1}(K\mathopen{(\!(}t\mathopen{)\!)},\Gamma)$, there exists a (unique) unramified $\Gamma$-torsor $P$ such that the twist $T\times^{\Gamma}P$ is strongly ramified, see [22, Lemmas 2.14, 2.16] for details.
Let $T$ a $\Gamma$-torsor over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$ and $\widetilde{T}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $T$. From the valuative criterion of properness, $\widetilde{T}$ is endowed with a $\Gamma$-action. We can form the stack quotient $[\widetilde{T}/\Gamma]$. Let $P$ be an unramified torsor such that $T\times^{\Gamma}P$ is strongly ramified. Since smooth base change commutes with normalization, we have $[\widetilde{T}\times^{\Gamma}P/\Gamma_{P}]\simeq[\widetilde{T}/\Gamma]$. Since the special fiber of $[\widetilde{T}\times^{\Gamma}P/\Gamma_{P}]$ is $BI_{T}$ and $I_{T}\cong\mu_{N}$, we get a morphism

### v4

@@ 1. Introduction
$O_{\gamma_{G}}=\int_{I_{\gamma_{G}}(F)\backslash G(F)}\mathbf{1}_{\mathfrak{g}(\mathcal{O}_{F})}(g^{-1}\gamma_{G}g)dg,$
@@
with $\gamma^{\prime}]$ running over the set of $G(F)$-conjugacy classes within the stable conjugacy class of $\gamma_{G}$. When $\kappa$ is the trivial character one gets the stable orbital integral $O^{\mathrm{stab}}_{\gamma_{G}}$. Thus, the problem of stabilization of orbital integrals can be reduced to expressing a $\kappa$-orbital integral for $G$ in terms of a stable orbital integral on some different unramified reductive group $H$ over $F$.
@@
The Hitchin fibration is endowed with an action of a Picard stack ${\mathcal{P}}_{G}\to\operatorname{\mathsf{A}}$. The anisotropic locus $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$ in $\operatorname{\mathsf{A}}$ is the locus of points $a$ such that every Higgs bundle $(E,\theta)\in\chi^{-1}(a)$ is generically anisotropic in the sense that for any trivialization of $E$ and $D$ over the generic point of $X$, the Higgs field $\theta$ is anisotropic as an element of $\mathfrak{g}(k(X))$. When $a$ belongs to $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, the group $\pi_{0}({\mathcal{P}}_{G,a})$ is finite. Fix $a\in\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, $H$ an endoscopic group of $G$, and $\kappa:\pi_{0}({\mathcal{P}}_{G,a})\to\overline{\mathbb{Q}_{\ell}}^{\times}$ the corresponding character. For technical reasons one needs to work on some étale open covering $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}\to\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$, and we denote by $\widetilde{\mathcal{M}}_{G}$, etc, the corresponding pullback. The locus $\widetilde{\operatorname{\mathsf{A}}}_{H}^{{\mathrm{ani}}}$ relative to $H$ is contained in $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}$.
@@
$^{\kappa}\,\widetilde{{\mathcal{M}}}_{G,a}(k)=q^{r_{G}^{H}(D)}\#^{\mathrm{\mathrm{stab}}}\,\widetilde{{\mathcal{M}}}_{H,a}(k).$
Here $^{\kappa}$ denotes the trace of Frobenius on the $\kappa$-isotypical component of the cohomology with compact supports, $^{\mathrm{stab}}$ denotes $^{\kappa}$ when $\kappa$ is the trivial character, $q$ is the cardinality of $k$ and $r_{G}^{H}(D)$ denotes $1/2(\mathrm{dim}(\widetilde{\mathcal{M}}_{G})-\mathrm{dim}(\widetilde{\mathcal{M}}_{H}))$.
@@ 2.1. Pseudo-finite fields
Let $\mathrm{\bf K}(\mathrm{Var}_{S})$ be the Grothendieck group of varieties over $S$ and $\mathcal{M}_{S}$ its localization by the class $\mathbb{L}$ of the affine line over $S$. Ivorra and Sebag prove in [31, Lemma 2.1] the existence of a unique ring morphism
@@
By additivity of $p_{!}$, as in the proof of [31, Lemma 2.1], this extends uniquely to a group morphism
@@
Property $(1)$ now follows from the standard fact that $\underline{\mathrm{Hom}}(V_{\alpha},M)^{G}\otimes V_{\alpha}$ is isomorphic to $M^{\alpha}$, see [11, 6.1] for details. Property $(2)$ follows from the fact that $M]=\sum_{\alpha\in G^{*}}[M^{\alpha}]$. We finally define $\chi_{S}(-,\alpha)$ for a virtual character $\alpha$ in the unique way such that $(3)$ holds. ∎
@@ 2.5. Denef-Pas language and definable subassignments
[…]
@@
Let $_{k}$ the final object in $\mathrm{GDef}_{\mathrm{Psf},k}$, whose value at every pseudo-finite extension of $k$ is the one point set. By definition
@@ 2.6. Evaluation of functions
Since $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k,x}$ is an expansion of $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k}$, by [9, Proposition 16.1.1], every $\psi\in\mathcal{C}_{\mathrm{Psf},k}(S)$ induces a unique function in $\mathcal{C}_{\mathrm{Psf},k,x}(S)$, that we still denote abusively $\psi$. By cell decomposition, or more precisely [9, Theorem 7.2.1], there exists an $\mathcal{L}_{\mathrm{DP},\mathrm{Psf},k,x}$-definable subassignment $V_{x}\subseteq S$, such that $x\in V_{x}(K)$, a definable function $g\colon V_{x}\to*$ and $\psi_{x}\in\mathcal{C}_{\mathrm{Psf},k,x}(*)$ such that $g^{*}(\psi_{x})=\psi$. We define the evaluation of $\psi$ at $(x,K)$ to be $\psi(x)=\psi_{x}\in\mathcal{C}_{\mathrm{Psf},k,x}(*)=\mathcal{C}_{\mathrm{Psf},k(x)}(*_{k(x)})$ .
@@ 2.7. Definable varieties
We shall usually consider such $\mathbf{X}$} as a definable in ${\mathbb{P}}^{n}$ with parameters $\pi({\mathbf{X}})$. In particular, by $\mathbf{X}$} is of pure dimension dd” we mean that the fibers ${\mathbf{X}}_{a}$ are points of varieties of pure dimension $d$.
@@
[…]
@@ 2.8. Fubini theorem
We say that $S\in\mathrm{GDef}_{\mathrm{Psf},k}$ is of dimension $d$ if $\mathrm{Kdim}(S)=d$, in the sense of [9, Section 3]. Roughly speaking, this means that the valued field part of $S$ is of dimension $d$. Fix $S\in\mathrm{Def}_{\mathrm{Psf},k}$ of dimension $d$ and let $\varphi\in\mathcal{C}_{\mathrm{Psf}}(S)$. Let $\left\lvert\omega_{0}\right\rvert_{S}$ be the canonical volume form, in the sense of [9, 15.1]. Let $C_{\mathrm{Psf}}^{d}(S)$ the quotient of $\mathcal{C}_{\mathrm{Psf}}(S)$ by the ideal of constructible functions with support included in a definable of dimension d1d-1. If the class $\varphi]\in C_{\mathrm{Psf}}^{d}(S)$ of $\varphi$ is integrable, then $\int_{S}[\varphi]\left\lvert\omega_{0}\right\rvert_{S}$ is defined in [9, 15.1] and belongs to $\mathcal{C}_{\mathrm{Psf}}(*_{k})$. Say that $\varphi\in\mathcal{C}_{\mathrm{Psf}}(S)$ is integrable if $\varphi]$ is integrable, and set
@@
Let $\mathbf{X}$} be a smooth definable algebraic variety over $k\mathopen{(\!(}t\mathopen{)\!)}$ of pure dimension $d$. Denote by $\left\lvert\omega_{\mathbf{X}}\right\rvert$ the definable volume form on $\mathbf{X}$} determined by the algebraic differential form of degree $d$ on ${\mathbf{X}}_{a}(K\mathopen{(\!(}t\mathopen{)\!)})$ for every pseudo-finite field. Let $\varphi\in\mathcal{C}_{\mathrm{Psf}}({\mathbf{X}})$. We say that $\varphi\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable if $\varphi]\left\lvert\omega_{\mathbf{X}}\right\rvert$ is integrable and set
@@ 2.9. Constructible functions with values in motives
Let $S\in\mathrm{GDef}_{\mathrm{Psf},k}$ be of dimension $d$. Let $\mathcal{C}^{\leq k}_{\mathrm{Psf}}(S)$ be the subring of $\mathcal{C}_{\mathrm{Psf}}(S)$ consisting of constructible functions with support contained in a definable subassignment of dimension at most $k$, and $C^{\leq k}_{\mathrm{Psf}}(S)=\mathcal{C}^{\leq k}_{\mathrm{Psf}}(S)/\mathcal{C}^{\leq{k-1}}_{\mathrm{Psf}}(S)$. Let $C_{\mathrm{Psf}}(S)=\bigoplus_{k\geq 0}C^{k}_{\mathrm{Psf}}(S)$. Recall from [9, Section 10] that the integral is constructed by defining for each morphism $f\colon X\to Y$ in $\mathrm{Def}_{\mathrm{Psf},k}$ a group homomorphism $f_{!}\colon IC_{\mathrm{Psf}}(X)\to C_{\mathrm{Psf}}(Y)$, where $IC_{\mathrm{Psf}}(X)$ is the subgroup of $C_{\mathrm{Psf}}(X)$ consisting of integrable functions. The integral of $\varphi]\in C_{\mathrm{Psf}}(X)$ is then defined as $f_{!}[\varphi]$, where $f\colon X\to*_{k}$ is the map to the final object. To prove Proposition 2.9.3, it is thus enough to prove that for every definable map $f\colon X\to Y$, $f_{!}\colon IC_{\mathrm{Psf}}(X)\to C_{\mathrm{Psf}}(Y)$ induces a map $IC_{\mathrm{mot}}(X)\to C_{\mathrm{mot}}(Y)$, where $IC_{\mathrm{mot}}(X)$ is the image of $IC_{\mathrm{Psf}}(X)$ by $\vartheta_{X}$.
@@
(2.9.4) ${IC_{\mathrm{Psf}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{!}}$$\scriptstyle{\vartheta_{S}}$${IC_{\mathrm{mot}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{!}}$${\mathcal{C}_{\mathrm{Psf}}(S^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\vartheta_{S^{\prime}}}$${\mathcal{C}_{\mathrm{mot}}(S^{\prime})}$
@@
[…]
@@ 2.13. Orbifold volume
A definable smooth Deligne-Mumford stack over $k\mathopen{[\![}t\mathopen{]\!]}$ is given by the following data: a definable groupoid $(V,A)$ such that $V/A$ is a definable algebraic variety $\mathbf{M}$} over $k\mathopen{[\![}t\mathopen{]\!]}$ together with a finite definable cover ${\mathbf{M}}=\cup{\mathbf{M}}_{i}$ such that the ${\mathbf{M}}_{i}$ form a Zariski cover of $\mathbf{M}$} into open pieces when specialized to a pseudo-finite field. We require as well that for all $i$, there is a smooth definable algebraic variety ${\mathbf{X}}_{i}$ and a definable finite algebraic group $\Gamma_{i}$ acting on ${\mathbf{X}}_{i}$ such that every orbit is contained in an affine subset, and such that the restriction of $(V,A)$ to ${\mathbf{M}}_{i}$ is identified with the quotient stack ${\mathbf{X}}_{i}/\Gamma_{i}]\to{\mathbf{X}}_{i}/\Gamma_{i}\simeq{\mathbf{M}}_{i}$, and there is a dense open subset $\mathbf{U}_{i}\subseteq{\mathbf{M}}_{i}$ over which the action is free.
@@
After shrinking the ${\mathbf{X}}_{i}$ if necessary, we may assume that their canonical bundles are trivial. We fix non-vanishing top degree forms $\omega_{i}$ on $X_{i}$, which induce, possibly pluricanonical, forms $\omega_{i,\mathrm{orb}}$ on ${\mathbf{M}}^{\natural}_{i}$ as in [23, Lemma 2.8] and the measures induced by $\omega_{i,\mathrm{orb}}$ glue to define a definable form $\left\lvert\omega_{\mathrm{orb}}\right\rvert$ on $\mathbf{M}$$^{\natural}$. We shall compute $\int_{{\mathbf{M}}^{\natural}}\left\lvert\omega_{\mathrm{orb}}\right\rvert$ in terms of the twisted inertia stack of $\mathcal{M}_{K}$.
@@
[…]
@@
Let $\widetilde{\mathcal{M}}_{x}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $\mathcal{M}_{x}$. By [23, Proposition 2.12], there is a finite totally ramified extension $L$ of $K\mathopen{(\!(}t\mathopen{)\!)}$ of degree $N$, with valuation ring $\mathcal{O}_{L}$ and residue field $k_{L}$, such that $\operatorname{\mathsf{Spec}}(\mathcal{O}_{L})/\mu_{N}]\simeq\widetilde{\mathcal{M}}_{x}$ and the induced isomorphism
$\operatorname{\mathsf{Spec}}(K)/\mu_{N}]=[\operatorname{\mathsf{Spec}}(k_{L})/\mu_{N}]\simeq\widetilde{\mathcal{M}}_{x,K}$
in the special fiber is independent of $L$ up to isomorphism. Hence we get a map $\operatorname{\mathsf{Spec}}(K)/\mu_{N}]\to\widetilde{\mathcal{M}}_{x,K}\to\mathcal{M}_{x,K}\to\mathcal{M}_{K}$, which gives a point $e(x)\in I_{\hat{\mu}}M_{K}(K)=I_{\hat{\mu}}{\mathbf{M}}(K)$.
@@
Before starting the proof, recall from [23, Section 2.5] some facts about $\Gamma$-torsors over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$, with $K$ pseudo-finite. A $\Gamma$-torsor $T=\operatorname{\mathsf{Spec}}(\prod_{i}L_{i})$ over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$ is said to be unramified if each $L_{i}$ is unramified. It is strongly ramified if each $L_{i}$ is totally ramified.
@@
For any $T$, from the description of $\mathrm{H}^{1}(K\mathopen{(\!(}t\mathopen{)\!)},\Gamma)$, there exists a (unique) unramified $\Gamma$-torsor $P$ such that the twist $T\times^{\Gamma}P$ is strongly ramified, see [23, Lemmas 2.14, 2.16] for details.
Let $T$ a $\Gamma$-torsor over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$ and $\widetilde{T}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $T$. From the valuative criterion of properness, $\widetilde{T}$ is endowed with a $\Gamma$-action. We can form the stack quotient $\widetilde{T}/\Gamma]$. Let $P$ be an unramified torsor such that $T\times^{\Gamma}P$ is strongly ramified. Since smooth base change commutes with normalization, we have $\widetilde{T}\times^{\Gamma}P/\Gamma_{P}]\simeq[\widetilde{T}/\Gamma]$. Since the special fiber of $\widetilde{T}\times^{\Gamma}P/\Gamma_{P}]$ is $BI_{T}$ and $I_{T}\cong\mu_{N}$, we get a morphism